OpenStudy (anonymous):

I NEED HELP!!!!!!! DETERMINE WHETHER THE TEST POINT IS A SOLUTION TO THE SYSTEM OF LINEAR INEQUALITIES ( 1, -2) x - 2 y > 3 2x + y < 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me?

OpenStudy (tkhunny):

You're not even going to show us that you tried? This is an exersize in three things, Substitution, Manipulation, and Notation. We have a point: (1,-2) The first thing we need is a Notation Test. The notation (1,-2) often means an ordered pair of x- and y-values. (1,-2) MEANS x = 1 and y = -2 We have two inequalities x - 2 y > 3 2x + y < 3 Now, it's time for substitution. We have x's and y's in these inequalities. We must subsitute the known values - known only if we understood the notation (1,-2) SUBSTITUTION x - 2 y > 3 2x + y < 3 or 1 - 2 (-2) > 3 2(1) + (-2) < 3 You must follow this, or we cannot get to the next step. MANIPULATION (a textbook might say, "after a little algebra") 1 - 2 (-2) > 3 2(1) + (-2) < 3 Resolve Parentheses 1 + 4 > 3 2 - 2 < 3 Addition 5 > 3 0 < 3 That's it. Now, we need to understand what we have done. We want to know if the point (1,-2) works in the two inequalities. After substituting the values, we have found an obvious result. 5 > 3 - which is true and 0 < 3 - which is true. If we understand the notation, "<" and ">", we will see that these raltionships are good and we are done with the problem. The given point DID prove appropriate for BOTH inequalities. The given point IS a solution to the SYSTEM of inequalities. Note: Had we achieved 0 > 3 or 10 < -5, we would have laughed at the result, because it is obviuosly false, and conclude that the given point is NOT a solution to the System. What do you think? Are we getting any closer? You will have to show some work if you are to get this in your head. I understand it takes a little courage.

OpenStudy (anonymous):

thank you i honestly didn't know how you gotten 1+4 > 3 but when you shown me step by step it made since because you 1 then you add 2+2=4 then that answer was 1+4> 3 2 (1) + (-2) < 3 then i said how did you get that then you explain it step by step 2 the 2- -2 I still dont know how you got the 2 -2 ? is it because there the same.

OpenStudy (tkhunny):

2 + (-2) = 2 - 2 2 - (-2) = 2 + 2 It is rather awkward notation. Youcan think about multiplicayion rules or direction on a number line. Taht may help, but you may just need to get used to it.